# Probing Loop Quantum Gravity via Kerr Black Hole and EHT Results

### v2

@@ IV Comparison With EHT Data
In this section, we will explore the constraints on the spin parameter $a$ for both quantum and Kerr BHs, using observational data from the EHT collaborations. These observations, focusing on M87* and Sgr A*, will allow us to determine the limits on the spin values of these BHs. We will compare the constraints obtained for both BHs to gain insights into how LQG might influence the characteristics of the rotating BH, providing a deeper understanding of the potential effects of quantum gravity on BHs. To do this, we calculate the angular radii of the shadows of both BHs to establish a comparative analysis with the angular radii of M87* and Sgr A*. Corresponding to such a bound on the spin parameter, the BH is considered to mimic either M87* or Sgr A* if the angular diameter of BH shadow falls within 1-$\sigma$ interval. This study specifically examines rotating BHs, as supermassive BHs are naturally expected to exhibit significant rotational features due to their formation and evolution processes. A coordinate-independent formalism, commonly referred to as the Kumar-Ghosh method [94, 95], is employed, where the shadow area is utilized which is defined as
@@
The impact of LQG that deviates the deformed rotating BH from Kerr BH is obvious in Fig. 4. We have plotted the shadow angular diameters of the Kerr and quantum Kerr BHs with respect to spin $a$ and compared it with M87* and Sgr A*. The 1-$\sigma$ uncertainty levels are indicated by dark gray regions. For the comparison with M87*, we found that the shadow diameter of Kerr BH lies within 1-$\sigma$ uncertainty level for $0\leq a\lesssim 0.586$, while, for other values of $a$, the shadow diameter of Kerr BH lies within 2-$\sigma$ uncertainty level. The upper bound of spin is denoted by $a_{\text{max}}\approx 0.586$ at which the transition of shadow diameter is observed from 1-$\sigma$ level to 2-$\sigma$. Therefore, the Kerr BH can be regarded identical with M87* for $0\leq a\lesssim 0.586$. However, the LQG BH does not mimic M87* as its shadow diameter lies within 2-$\sigma$ uncertainty level for all values of spin which is a tremendous impact of LQG on quantum Kerr BH. Generally, the 2-$\sigma$ uncertainty level is also considered for the comparison, however, we ignore this region in order to reduce the possibility of uncertainty. This essentially eliminates the possibility for the LQG BH to behave like M87*. However, an entirely opposite impact of LQG is visualized for the case of Sgr A*. The Kerr BH behaves identical with Sgr A* for all values of spin as its shadow diameter lies within 1-$\sigma$ uncertainty level and $\theta_{\text{d}}$ approaches the median value $48.7\mu\text{as}$ as $a$ approaches its maximal value. For the LQG BH, the shadow diameter also lies within 1-$\sigma$ uncertainty level. However, under the influence of LQG, the shadow diameter for LQG BH is closer to the median value $48.7\mu\text{as}$ for each value of spin. Therefore, one may regard the LQG BH to mimic Sgr A* for all values of spin which is more likely than the Kerr BH within its interval of spin.
@@ V Effect of Plasma on Shadow
We know that astrophysical BHs are generally surrounded by plasma. Moreover, excited states of matter at high temperatures with their constituent particles (positive and negative charges) may have significant impact on the LQG BH at Planck scales. This being an important aspect in theoretical physics, we investigate the influence of plasma on the appearance of quantum Kerr BH. We consider pressureless and non-magnetized plasma with distribution functions dependent on $r$ and $\theta$. We begin by considering the plasma electron frequency defined as
@@
Figure 5: Influence of $\omega_{\text{c}}$ in case I (upper panel) and case II (lower panel) on shadows of quantum Kerr BH for a fixed value of $a$.
@@
For both of above mentioned cases, we have plotted the shadows in Fig. 5 for a few values of plasma parameter $\omega_{\text{c}}$ and keeping a fixed spin of the BH. The spin deviates the shadow from a pure circular loop causing a distortion. While, the plasma parameter $\omega_{\text{c}}$ reduces the size of the shadow. It can also be seen that $\omega_{\text{c}}$ in case II is more significant and has higher sensitivity as compared to the case I. The variation in the shadow size for case II is greater than the variation observed in case I.
@@ VI Discussion
By a smooth matching of APS and qOS spacetimes by identifying a set of suitable coordinate transformation, a deformed Schwarzschild BH in LQG is obtained with a relatively weak extra term of order $r^{-4}$ depending on Barbero-Immirzi parameter $\gamma$ [28]. During the collapse of the dust ball, corresponding to a lower bound of the radial coordinate, a lower bound on the mass of the quantum Kerr BH is obtained. Below which, there exists no horizon for the LQG BH, whereas, a two horizon system is obtained for all mass values above the minimum mass limit. Ye et al. [29] studied the shadows of static LQG BH to determine the effect of LQG on the Schwarzschild BH. They considered the fixed value of the Barbero-Immirzi parameter and hence the parameter $\alpha$. They presented their findings in comparison with the results of Schwarzschild BH. Motivated by this, we considered the rotating counterpart of the deformed Schwarzschild BH in LQG and accomplished our analysis in comparison with the results for Kerr BH.
@@
The rotating BH metric (7) encompasses various features of Kerr-like BH, especially the existence of time translation and rotational invariance isometries. Additionally, on removing the spin, its exact static counterpart is recovered, and on removing the quantum effects, the rotating metric reduces to Kerr metric and the static metric reduces to Schwarzschild metric. The parameter $\alpha$ though appears as the LQG parameter in the BH metric, however, it does not behave as a free parameter and has a fixed value that cannot be varied. The metric function of quantum Kerr BH is expressed in terms of the metric function of Kerr BH for which the effect of LQG can be observed in the Eq. (10) and from the Fig. 1. The Kerr BH in LQG also exhibits two horizons, though the event horizon being small as compared to the event horizon of Kerr BH, reduces the size of BH due to LQG effects. Its extreme spin value is also reduced to the half of extreme spin of Kerr BH.
@@
The function $\mathcal{R}(r)$ plays a vital role in particle orbits and shadow analysis. Like the metric function, it is also expressed in terms of the function corresponding to the Kerr BH which generates an identical form of effective potential as well. This form enables us to generalize the location of unstable null orbits for such kind of metrics. We proved a theorem based on convexity and the familiar results of effective potential of Kerr BH, which ensures that the unstable orbits for such a deformed Kerr BH will be smaller than the unstable orbits for Kerr BH. This result holds true only if the extra term in effective potential function is decreasing function of $r$. The result in this theorem is then verified numerically in the Fig. 2. The impact parameters and celestial coordinates for quantum Kerr BH are also expressed in terms of the impact parameters and celestial coordinates for Kerr BH. The deviation terms in the celestial coordinates determine the deviation of shadow of quantum Kerr BH from the shadow of Kerr BH. The LQG effects influences both the size of the shadow and the distortion in it. Since, the extreme spin for quantum Kerr BH is $\sim$0.4952, so the ergosphere does not get enough strength to create a flat shadow like the shadow of Kerr BH.
@@
The astrophysical impact of LQG on quantum Kerr BH is investigated through the comparison of the shadow size for Kerr and quantum Kerr BHs with the size of M87* and Sgr A*. Along with the images, the data from EHT results enabled us to draw an analysis to determine the constraints on the BH spin parameters and the influence of LQG on it. The Kerr BH is defined for $0\leq a\leq 1$, however, it becomes identical with M87* for $0\leq a\leq 0.586$. Whereas, the effect of LQG on quantum Kerr BH inhibits it to behave like M87* for all spin values. On the other hand, both Kerr and quantum Kerr BHs mimic Sgr A* for all spin values, however, due to LQG effects, the quantum Kerr BH is more likely to mimic Sgr A* than Kerr BH.
@@
The plasma surrounding the LQG BH has a great impact on the light propagation. The shadow size is reduced by increasing the value of $\omega_{\text{c}}$ in both cases. However, the quantity of variation in shadow size reveals that $\omega_{\text{c}}$ in case II is more sensitive and has greater impact than that in case I.
@@
As a future project, one may investigate the effect of LQG on the deflection of light in strong and weak regimes of quantum Kerr BH. The study of BH evaporation rate, Hawking radiation via tunneling process and Unruh effect would also be an intriguing analysis. Moreover, it will be interesting to study the effect of LQG on the quantum Kerr BH behaving as a particle accelerator.
@@ Acknowledgment
This paper is dedicated to the memory of Prof. Jerzy Lewandowski, who recently derived the original deformed BH metric in LQG which is considered in this paper. He is no more with us but his work will continue to inspire us, leaving a lasting legacy in theoretical physics.

### v3

@@ IV Comparison With EHT Data
In this section, we will explore the constraints on the spin parameter $a$ for both Kerr and quantum corrected Kerr BHs, using observational data from the EHT collaborations. These observations, focusing on M87* and Sgr A*, will allow us to determine the limits on the spin values of these BHs. We will compare the constraints obtained for both BHs to gain insights into how LQG might influence the characteristics of the rotating BH, providing a deeper understanding of the potential effects of quantum gravity on BHs. To do this, we calculate the angular radii of the shadows of both BHs to establish a comparative analysis with the angular radii of M87* and Sgr A*. Corresponding to such a bound on the spin parameter, the BH is considered to mimic either M87* or Sgr A* if the angular diameter of BH shadow falls within 1-$\sigma$ interval. This study specifically examines rotating BHs, as supermassive BHs are naturally expected to exhibit significant rotational features due to their formation and evolution processes. A coordinate-independent formalism, commonly referred to as the Kumar-Ghosh method [93, 94], is employed, where the shadow area is utilized which is defined as
@@
The impact of LQG that deviates the quantum corrected Kerr BH from Kerr BH is obvious in Fig. 4. We have plotted the shadow angular diameters of the Kerr and quantum corrected Kerr BHs with respect to spin $a$ and compared it with M87* and Sgr A*. The 1-$\sigma$ uncertainty levels are indicated by dark gray regions. For the comparison with M87*, we found that the shadow diameter of Kerr BH lies within 1-$\sigma$ uncertainty level for $0\leq a\lesssim 0.586$, while, for other values of $a$, the shadow diameter of Kerr BH lies within 2-$\sigma$ uncertainty level. The upper bound of spin is denoted by $a_{\text{max}}\approx 0.586$ at which the transition of shadow diameter is observed from 1-$\sigma$ level to 2-$\sigma$. Therefore, the Kerr BH can be regarded identical with M87* for $0\leq a\lesssim 0.586$. However, the quantum corrected Kerr BH does not mimic M87* as its shadow diameter lies within 2-$\sigma$ uncertainty level for all values of spin which is a tremendous impact of LQG on quantum corrected Kerr BH. Generally, the 2-$\sigma$ uncertainty level is also considered for the comparison, however, we ignore this region in order to reduce the possibility of uncertainty. This essentially eliminates the possibility for the quantum corrected Kerr BH to behave like M87*. However, an entirely opposite impact of LQG is visualized for the case of Sgr A*. The Kerr BH behaves identical with Sgr A* for all values of spin as its shadow diameter lies within 1-$\sigma$ uncertainty level and $\theta_{\text{d}}$ approaches the median value $48.7\mu\text{as}$ as $a$ approaches its maximal value. For the quantum corrected Kerr BH, the shadow diameter also lies within 1-$\sigma$ uncertainty level. However, under the influence of LQG, the shadow diameter for quantum corrected Kerr BH is closer to the median value $48.7\mu\text{as}$ for each value of spin. Therefore, one may regard the quantum corrected Kerr BH to mimic Sgr A* for all values of spin which is more likely than the Kerr BH within its interval of spin.
@@ V Effect of Plasma on Shadow
We know that astrophysical BHs are generally surrounded by plasma. Moreover, excited states of matter at high temperatures with their constituent particles (positive and negative charges) may have significant impact on the quantum corrected Kerr BH at Planck scales. This being an important aspect in theoretical physics, we investigate the influence of plasma on the appearance of quantum corrected Kerr BH. We consider pressureless and non-magnetized plasma with distribution functions dependent on $r$ and $\theta$. We begin by considering the plasma electron frequency defined as
@@
Figure 5: Influence of $\omega_{\text{c}}$ in case I (upper panel) and case II (lower panel) on shadows of quantum corrected Kerr BH for a fixed value of $a$.
@@
For both of above mentioned cases, we have plotted the shadows in Fig. 5 for a few values of plasma parameter $\omega_{\text{c}}$ and keeping a fixed spin of the quantum corrected Kerr BH. The spin deviates the shadow from a pure circular loop causing a distortion. While, the plasma parameter $\omega_{\text{c}}$ reduces the size of the shadow. It can also be seen that $\omega_{\text{c}}$ in case II is more significant and has higher sensitivity as compared to the case I. The variation in the shadow size for case II is greater than the variation observed in case I.
@@ VI Discussion
By a smooth matching of APS and qOS spacetimes by identifying a set of suitable coordinate transformation, a quantum corrected Schwarzschild BH in LQG is obtained with a relatively weak extra term of order $r^{-4}$ depending on Barbero-Immirzi parameter $\gamma$ [28]. During the collapse of the dust ball, corresponding to a lower bound of the radial coordinate, a lower bound on the mass of this BH is obtained, below which, there exists no horizon, whereas, a two horizon system is obtained for all mass values above the minimum mass limit. Ye et al. [29] studied the shadows of this BH to determine the effect of LQG on the BH. They considered the fixed value of the Barbero-Immirzi parameter and hence the parameter $\alpha$. They presented their findings in comparison with the results of Schwarzschild BH. Motivated by this, we considered the rotating counterpart of the quantum corrected Schwarzschild BH, in effective LQG and accomplished our analysis in comparison with the results for Kerr BH.
@@
The rotating BH metric (7) being an effective metric still encompasses various features of Kerr-like BH, especially the existence of time translation and rotational invariance isometries. Additionally, on removing the spin, its exact static counterpart is recovered, and on removing the quantum effects, the rotating metric reduces to Kerr metric and the static metric reduces to Schwarzschild metric. The parameter $\alpha$ though appears as the LQG parameter in the BH metric, however, it does not behave as a free parameter and has a fixed value that cannot be varied. The metric function of the quantum corrected Kerr BH is expressed in terms of the metric function of Kerr BH for which the effect of LQG can be observed in the Eq. (10) and from the Fig. 1. The quantum corrected Kerr BH in effective LQG also exhibits two horizons, though the event horizon being small as compared to the event horizon of Kerr BH, reduces the size of BH due to LQG effects. Its extreme spin value is also reduced to the half of extreme spin of Kerr BH.
@@
The function $\mathcal{R}(r)$ plays a vital role in particle orbits and shadow analysis. Like the metric function, it is also expressed in terms of the function corresponding to the Kerr BH which generates an identical form of effective potential as well. This form enables us to generalize the location of unstable null orbits for such kind of metrics. We proved a theorem based on convexity and the familiar results of effective potential of Kerr BH, which ensures that the unstable orbits for such a quantum corrected Kerr BH will be smaller than the unstable orbits for Kerr BH. This result holds true only if the extra term in effective potential function is decreasing function of $r$. The result in this theorem is then verified numerically in the Fig. 2. The impact parameters and celestial coordinates for quantum corrected Kerr BH are also expressed in terms of the impact parameters and celestial coordinates for Kerr BH. The deviation terms in the celestial coordinates determine the deviation of shadow of quantum corrected Kerr BH from the shadow of Kerr BH. The LQG effects influences both the size of the shadow and the distortion in it. Since, the extreme spin for quantum corrected Kerr BH is $\sim$0.4952, so the ergosphere does not get enough strength to create a flat shadow like the shadow of Kerr BH.
@@
The astrophysical impact of LQG on quantum corrected Kerr BH is investigated through the comparison of the shadow size for Kerr and quantum corrected Kerr BHs with the size of M87* and Sgr A*. Along with the images, the data from EHT results enabled us to draw an analysis to determine the constraints on the BH spin parameters and the influence of LQG on it. The Kerr BH is defined for $0\leq a\leq 1$, however, it becomes identical with M87* for $0\leq a\leq 0.586$. Whereas, the effect of LQG on quantum corrected Kerr BH inhibits it to behave like M87* for all spin values. On the other hand, both Kerr and quantum corrected Kerr BHs mimic Sgr A* for all spin values, however, due to LQG effects, the quantum corrected Kerr BH is more likely to mimic Sgr A* than Kerr BH.
@@
The plasma surrounding the quantum corrected Kerr BH in effective LQG has a great impact on the light propagation. The shadow size is reduced by increasing the value of $\omega_{\text{c}}$ in both cases. However, the quantity of variation in shadow size reveals that $\omega_{\text{c}}$ in case II is more sensitive and has greater impact than that in case I.
@@
As a future project, one may investigate the effect of LQG on the deflection of light in strong and weak regimes of quantum corrected Kerr BH. The study of BH evaporation rate, Hawking radiation via tunneling process and Unruh effect would also be an intriguing analysis. Moreover, it will be interesting to study the effect of LQG on the quantum corrected Kerr BH behaving as a particle accelerator.
@@ Acknowledgment
This paper is dedicated to the memory of Prof. Jerzy Lewandowski, who derived the static quantum corrected BH metric in LQG which is considered in this paper. He is no more with us but his work will continue to inspire us, leaving a lasting legacy in theoretical physics. We are thankful to Francesco Fazzini for discussions and useful suggestions for improvements in the paper.